# Fixed-sparsity matrix approximation from matrix-vector products

### v1

@@ Abstract
We study the problem of approximating a matrix $\mathbf{A}$ with a matrix that has a fixed sparsity pattern (e.g., diagonal, banded, etc.), when $\mathbf{A}$ is accessed only by matrix-vector products. We describe a simple randomized algorithm that returns an approximation with the given sparsity pattern with Frobenius-norm error at most $(1+\varepsilon)$ times the best possible error. When each row of the desired sparsity pattern has at most $s$ nonzero entries, this algorithm requires $O(s/\varepsilon)$ non-adaptive matrix-vector products with $\mathbf{A}$. We proceed to prove a matching lower-bound. Specifically, we show that for any $s\geq 1$, there are matrices $\mathbf{A}$ such that, for any sparsity pattern with $\Theta(s)$ nonzeros per row and column, any algorithm which obtains a $(1+\varepsilon)$ accurate approximation of the given sparsity from matrix-vector products requires at least $\Omega(s/\varepsilon)$ matrix-vector products. Our bounds therefore resolve the matrix-vector product query complexity of the problem up to constant factors, even for the well-studied case of diagonal approximation, for which no previous lower bounds were known.
@@ 1 Introduction
Learning about a matrix $\mathbf{A}$ from matrix-vector product (matvec) queries [^1] In some settings, it may also be possible to do matrix-transpose-vector queries $\mathbf{y}\mapsto\mathbf{A}^{\mathsf{T}}\mathbf{y}$. It is an open question to understand when matrix-transpose-vector queries may be beneficial [BHOT24]. $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\mapsto\mathbf{A}\mathbf{x}_{1},\ldots,\mathbf{A}\mathbf{x}_{m}$ is a widespread task in numerical linear algebra, theoretical computer science, and machine learning. Algorithms that only access $\mathbf{A}$ through matvec queries (also known as matrix-free algorithms) are useful when $\mathbf{A}$ is not known explicitly, but admits efficient matvecs. Common settings include when $\mathbf{A}$ is a solution operator for a linear partial differential equation [Kar+21, SO21, BET22], a function of another matrix that can be applied with Krylov subspace methods [GS92, Hig08], the Hessian of a neural network that can be applied via back-propagation [Pea94], or a data matrix in compressed sensing applications [KK11]. Matrix-free algorithms are often the methods of choice due to practical considerations such as memory usage and data movement. If the queries are chosen non-adaptively — i.e, the $i$-th query vector $\mathbf{x}_{i}$ does not depend on the results $\mathbf{Ax}_{1},\ldots,\mathbf{Ax}_{i-1}$ from previous queries — queries can be parallelized, potentially leading to substantial runtime improvements [Mur+23]. Non-adaptive queries can also be computed in a single pass over $\mathbf{A}$, and thus, non-adaptive matvec query algorithms are an important tool in streaming and distributed computation [CW09, MT20].
@@
In many cases, the cost of matrix-free algorithms is dominated by the cost of the matvec queries. As such, a key goal is to understand the minimum number of queries required to solve a given problem, also known as the query complexity. Any algorithm automatically yields an upper bound on the query complexity, whereas it can be more challenging to prove lower bounds. Problems for which the matvec query complexity have been extensively studied include low-rank approximation [HMT11, SER18, TW23, BCW22, BN23, HT23], spectrum approximation [Ski89, WWAF06, SW23], trace and diagonal estimation [Gir87, Hut89, BKS07, MMMW21], the approximation of linear system solutions and the action of matrix functions [GS92, Gre97, BHSW20], and matrix property testing [SWYZ21, NSW22].
@@
An important class of problems considers recovering $\mathbf{A}$ from matvec queries when $\mathbf{A}$ is structured, or relatedly, finding the nearest approximation to $\mathbf{A}$ within a structured class of matrices. Example classes that have been studied extensively in this setting include low-rank matrices [HMT11, TW23], hierarchical low-rank matrices [LLY11, Mar16, LM22, LM22a, SO21, HT23], diagonal matrices [BKS07, TS11, BN22, DM23], sparse matrices [CPR74, CM83, CC86, WEV13, DSBN15], and beyond [WSB11, SKO21].
@@ 1.1 Fixed-sparsity matrix approximation
Hence, Problem 1 is asking for a near-optimal approximation to $\mathbf{A}$ of the given sparsity $\mathbf{S}$ [^2] This is reminiscent of the low-rank approximation problem, in which we aim to find a rank-$k$ approximation to $\mathbf{A}$ competitive with the best rank-$k$ approximation [HMT11, TW23]. In particular, if $\mathbf{A}$ already has sparsity pattern $\mathbf{S}$, then $\mathbf{A}=\mathbf{S}\circ\mathbf{A}$, so solving Problem 1 will recover $\mathbf{A}$ exactly.
@@ 1.2 Our Contributions and Roadmap
Thus, $\Omega(s/\varepsilon)$ queries are required to solve Problem 1 with any reasonable probability. Note that since solving Problem 1 gives a solution to Problem 2, the lower bound Theorem 2 also implies that $\Omega(s/\varepsilon)$ queries are required to solve Problem 2. Our approach uses an invariance property of Wishart matrices after a sequence of adaptive queries [BHSW20]. Note that since our hard instance is symmetric, Theorem 2 also holds against algorithms which are allowed to use matvec queries with $\mathbf{A}^{\mathsf{T}}$.
Our lower bound applies even to the well-studied case of best approximation by a diagonal matrix, and more broadly, to approximation by a banded matrix. To the best of our knowledge, our lower bounds are the first (adaptive or non-adaptive) for Problem 1 (and thus for Problem 2), even for the well-studied case of diagonal approximation.
In Section 4, we compare our algorithm to widely-used coloring-based methods for fixed-sparsity pattern matrix approximation [CPR74, CM83, CC86, etc.]. We show that there are situations where coloring methods perform worse than the algorithm described in this paper by a quadratic factor or more. Specifically, even for matrices with $\leq s$ non-zeros per row and column, they can require $O(s^{2})$ instead of $O(s)$ matvec queries. We also discuss a setting in which coloring methods can outperform the algorithm from this paper.
@@ 1.3 Past work
To the best of our knowledge, Problems 1 and 2 have not be previously stated explicitly in the given generality. However, there is a range of past work that studies special cases of these problems. We categorize these works into the zero error case, where $\mathbf{A}=\mathbf{S}\circ\mathbf{A}$, and the nonzero error case, where $\mathbf{A}\neq\mathbf{S}\circ\mathbf{A}$. In addition, we discuss how Problems 1 and 2 differ from the standard sparse-recovery problem in compressed sensing.
@@ 1.3.1 Zero error
Some of the earliest work relating to Problems 1 and 2 seeks to recover Jacobian and Hessian matrices with a known sparsity pattern from matvecs [CPR74, CM83, CC86]. These methods make use of the fact that a graph coloring of a particular graph, induced by the sparsity pattern of $\mathbf{A}$, can be used to obtain a set of query vectors which are sufficient to exactly recover $\mathbf{A}$. In many (but not all) cases, exact recovery is possible with $s$ queries. Coloring methods have also influenced many algorithms for the nonzero error setting. We compare our results to such coloring-based methods in Section 4, arguing that Algorithm 1 always performs better in the zero-error setting, since it always requires just $s$ queries.
More generally, [HT23] studies the matvec query complexity of exact recovery of a wide range of linearly parameterized matrix families, proving matching upper and lower-bounds on the number of queries required. In particular, it is shown that recovering a diagonal matrix requires one query, recovering a block-diagonal matrix with $s\times s$ blocks requires $s$ queries, and recovering a tridiagonal matrix requires $3$ queries. Recovering a general $n\times d$ matrix is shown to require $d$ queries. With probability one, Algorithm 1 matches these lower bounds (see Proposition 1).
@@ 1.3.2 Nonzero error
$\mathbf{d}_{m}=\big{[}\sum_{j=1}^{m}\mathbf{r}_{j}\circ(\mathbf{A}\mathbf{r}_{j})\Big{]}\ooalign{\hss$\circ$\cr$\div$}\Big{[}\sum_{j=1}^{m}\mathbf{r}_{j}\circ\mathbf{r}_{j}\Big{]},$ (3)
@@
where “ $\circ$ $\div$ ” indicates entrywise division and the entries of the vectors $\mathbf{r}_{j}$ are all independent random variables with mean zero and variance one. A number of analyses of this estimator have been given for various distributions [BKS07, TS11, BN22, HIS23, DM23]. In particular, [BN22, DM23] give error bounds for Problem 2, showing it suffices to set $m=O(1/\varepsilon)$, matching Theorem 1 up to constant factors. In fact, when $\mathbf{S}=\mathbf{I}$ our Algorithm 1 is equivalent to Equation 3 if the query vectors $\mathbf{r}_{i}$ are Gaussian. We detail this connection in Remark 3.
@@
Past work has also studied Problem 1 with the goal of approximating a potentially non-sparse matrix by a sparse matrix. For instance, there is a long line of work on approximating matrix functions $\mathbf{A}=f(\mathbf{H})$ from matvecs. If $\mathbf{H}$ is sparse, then the entries of $\mathbf{A}=f(\mathbf{H})$ decay exponentially away from the nonzero entries of $\mathbf{H}$ under mild assumptions on $f(x)$ [DMS84, BR08, BBR13]. As such, it is reasonable to approximate $\mathbf{A}$ with a sparse matrix of a sparsity similar to $\mathbf{H}$. This observation has been used in matrix approximation algorithms [TS11, SLO13, FSS21, PN23]. Broadly speaking, these algorithms aim to combine the coloring methods described above with the estimator $\mathbf{d}_{m}$ described in Equation 3. For banded matrices, one can use the existing analyses of $\mathbf{d}_{m}$ to analyze the performance of these methods, showing that they solve Problem 2 to accuracy ${\varepsilon}$ using $O(s/\varepsilon)$ matrix-vector queries. We include a note on this in Section 4.1, as we were unable to find such an analysis in the literature. Theorem 1 shows that Algorithm 1 matches this bound for arbitrary sparsity patterns.
@@
More recently, motivated by the field of partial differential equation (PDE) learning, there has been widespread interest in learning the solution operators of PDEs from input-output data of forcing terms and solutions, analogous to matrix-vector products [SKO21, SO21, BHT23, Kar+21]. The method in [SO21] obtains a fixed-sparsity approximation to the sparse Cholesky factorization of the solution operator by coloring, which is provably accurate for certain problems. This makes use of the fact that in certain settings a fixed-sparsity Cholesky factorization is accurate and can be efficiently computed [SKO21]. This is broadly related to the (factorized) sparse approximate inverse problem for obtaining preconditioners [BT99]. The method in [BHT23] also derives a continuous analogue of a generalized coloring algorithm for targeting low-rank subblocks of hierarchical matrices [LM22a], and then recovering these subblocks using the randomized SVD [HMT11]. The final step of this algorithm reduces to the recovery of a block diagonal matrix.
@@ 1.3.3 The sparse recovery problem in compressed-sensing
It is important to contrast the aims and methods of this paper with the rich literature on compressed sensing and sparse recovery [EK12, FR13, etc.].
@@
Critically, the support of $\widetilde{\mathbf{a}}$ is not known ahead of time. In fact, if the support were specified, this would be a trivial problem; simply take $\mathbf{M}$ to have $k$ rows, each a standard basis vector corresponding to an entry of the support.
A number of past works [WSB11, WEV13, DSBN15, etc.] have also studied a matrix version of this problem in which one aims to obtain an $sd$-sparse matrix $\widetilde{\mathbf{A}}$ for which
@@
using only bi-linear measurements of $\mathbf{A}$: $(\mathbf{U},\mathbf{V})\mapsto\mathbf{U}\mathbf{A}\mathbf{V}^{\mathsf{T}}$. Note that the matrix problem is actually equivalent to a restricted version of the vector problem. Indeed, if $\mathbf{a}=\operatorname{vec}(\mathbf{A})$ and $\mathbf{M}=\mathbf{U}\otimes\mathbf{V}$, then $\mathbf{M}\mathbf{a}=\operatorname{vec}(\mathbf{U}\mathbf{A}\mathbf{V}^{\mathsf{T}})$. Here $\operatorname{vec}(\cdot)$ forms a vector by stacking the columns of a matrix on top of one another and $\otimes$ is the Kronecker product.
Even for the general vector recovery problem, such algorithms necessarily have worse dependencies on $\varepsilon$ and $d$ than the bounds we prove for the algorithm described in the next section. In particular, algorithms solving the $\ell_{2}/\ell_{2}$ sparse recovery problem necessarily require $\Omega(s/\varepsilon^{2}+s\log(ds)/\varepsilon)$ linear measurements [PW11]. In fact, even if the problem is relaxed, so that the output vector $\widetilde{\mathbf{a}}$ is allowed to be non-sparse, $\Omega(s\log(d/s)/\varepsilon)$ queries are required [PW11]. In contrast, our upper-bound Theorem 1/Corollary 1 have no dependence on the dimension $d$.

### v3

@@ Abstract
We study the problem of approximating a matrix $\mathbf{A}$ with a matrix that has a fixed sparsity pattern (e.g., diagonal, banded, etc.), when $\mathbf{A}$ is accessed only by matrix-vector products. We describe a simple randomized algorithm that returns an approximation with the given sparsity pattern with Frobenius-norm error at most $(1+\varepsilon)$ times the best possible error. When each row of the desired sparsity pattern has at most $s$ nonzero entries, this algorithm requires $O(s/\varepsilon)$ non-adaptive matrix-vector products with $\mathbf{A}$. We also prove a matching lower-bound, showing that, for any sparsity pattern with $\Theta(s)$ nonzeros per row and column, any algorithm achieving $(1+\epsilon)$ approximation requires $\Omega(s/\varepsilon)$ matrix-vector products in the worst case. We thus resolve the matrix-vector product query complexity of the problem up to constant factors, even for the well-studied case of diagonal approximation, for which no previous lower bounds were known.
@@ 1 Introduction
Learning about a matrix $\mathbf{A}$ from matrix-vector product (matvec) queries [^1] In some settings, it may also be possible to do matrix-transpose-vector queries $\mathbf{y}\mapsto\mathbf{A}^{\mathsf{T}}\mathbf{y}$. It is an open question to understand when matrix-transpose-vector queries may be beneficial [boulle_halikias_otto_townsend_24]. $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\mapsto\mathbf{A}\mathbf{x}_{1},\ldots,\mathbf{A}\mathbf{x}_{m}$ is a widespread task in numerical linear algebra, theoretical computer science, and machine learning. Algorithms that only access $\mathbf{A}$ through matvec queries (also known as matrix-free algorithms) are useful when $\mathbf{A}$ is not known explicitly, but admits efficient matvecs. Common settings include when $\mathbf{A}$ is a solution operator for a linear partial differential equation [karniadakis_kevrekidis_lu_perdikaris_wang_yang_21, schafer_owhadi_21, boulle_earls_townsend_22], a function of another matrix that can be applied with Krylov subspace methods [gallopoulos_saad_92, higham_08], the Hessian of a neural network that can be applied via back-propagation [pearlmutter_94], or a data matrix in compressed sensing applications [kalender_kalender_11]. Matrix-free algorithms are often the methods of choice due to practical considerations such as memory usage and data movement. If the queries are chosen non-adaptively — i.e, the $i$-th query vector $\mathbf{x}_{i}$ does not depend on the results $\mathbf{Ax}_{1},\ldots,\mathbf{Ax}_{i-1}$ from previous queries — queries can be parallelized, potentially leading to substantial runtime improvements [murray_demmel_mahoney_erichson_melnichenko_malik_grigori_luszczek_derezinski_lopes_liang_luo_dongarra_23]. Non-adaptive queries can also be computed in a single pass over $\mathbf{A}$, and thus, non-adaptive matvec query algorithms are an important tool in streaming and distributed computation [clarkson_woodruff_09, martinsson_tropp_20].
@@
In many cases, the cost of matrix-free algorithms is dominated by the cost of the matvec queries. As such, a key goal is to understand the minimum number of queries required to solve a given problem, also known as the query complexity. Any algorithm automatically yields an upper bound on the query complexity, whereas it can be more challenging to prove lower bounds. Problems for which the matvec query complexity have been extensively studied include low-rank approximation [halko_martinsson_tropp_11, simchowitz_elalaoui_recht_18, tropp_webber_23, bakshi_clarkson_woodruff_22, bakshi_narayanan_23, halikias_townsend_23], spectrum approximation [skilling_89, weisse_wellein_alvermann_fehske_06, swartworth_woodruff_23], trace and diagonal estimation [girard_87, hutchinson_89, bekas_kokiopoulou_saad_07, meyer_musco_musco_woodruff_21], the approximation of linear system solutions and the action of matrix functions [gallopoulos_saad_92, greenbaum_97, braverman_hazan_simchowitz_woodworth_20], and matrix property testing [sun_woodruff_yang_zhang_21, needell_swartworth_woodruff_22].
@@
An important class of problems considers recovering $\mathbf{A}$ from matvec queries when $\mathbf{A}$ is structured, or relatedly, finding the nearest approximation to $\mathbf{A}$ within a structured class of matrices. Example classes that have been studied extensively in this setting include low-rank matrices [halko_martinsson_tropp_11, tropp_webber_23], hierarchical low-rank matrices [lin_lu_ying_11, martinsson_16, levitt_martinsson_22, levitt_martinsson_22a, schafer_owhadi_21, halikias_townsend_23], diagonal matrices [bekas_kokiopoulou_saad_07, tang_saad_11, baston_nakatsukasa_22, dharangutte_musco_23], sparse matrices [curtis_powell_reid_74, coleman_more_83, coleman_cai_86, wimalajeewa_eldar_varshney_13, dasarathy_shah_bhaskar_nowak_15], and beyond [waters_sankaranarayanan_baraniuk_11, schafer_katzfuss_owhadi_21].
@@ 1.1 Fixed-sparsity matrix approximation
Hence, Problem 1 is asking for a near-optimal approximation to $\mathbf{A}$ of the given sparsity $\mathbf{S}$ [^2] This is reminiscent of the low-rank approximation problem, in which we aim to find a rank-$k$ approximation to $\mathbf{A}$ competitive with the best rank-$k$ approximation [halko_martinsson_tropp_11, tropp_webber_23]. In particular, if $\mathbf{A}$ already has sparsity pattern $\mathbf{S}$, then $\mathbf{A}=\mathbf{S}\circ\mathbf{A}$, so solving Problem 1 will recover $\mathbf{A}$ exactly.
@@ 1.2 Our Contributions and Roadmap
Thus, $\Omega(s/\varepsilon)$ queries are required to solve Problem 1 with any reasonable probability. Note that since solving Problem 1 gives a solution to Problem 2, the lower bound Theorem 2 also implies that $\Omega(s/\varepsilon)$ queries are required to solve Problem 2. Our approach uses an invariance property of Wishart matrices after a sequence of adaptive queries [braverman_hazan_simchowitz_woodworth_20]. Note that since our hard instance is symmetric, Theorem 2 also holds against algorithms which are allowed to use matvec queries with $\mathbf{A}^{\mathsf{T}}$.
Our lower bound applies even to the well-studied case of best approximation by a diagonal matrix, and more broadly, to approximation by a banded matrix. To the best of our knowledge, our lower bounds are the first (adaptive or non-adaptive) for Problems 1 and 2 even for these special cases.
In Section 4, we compare our algorithm to widely-used coloring-based methods for fixed-sparsity pattern matrix approximation [curtis_powell_reid_74, coleman_more_83, coleman_cai_86, etc.]. We show that there are situations where coloring methods perform worse than the algorithm described in this paper by a quadratic factor or more. Specifically, even for matrices with $\leq s$ non-zeros per row and column, they can require $O(s^{2})$ instead of $O(s)$ matvec queries. We also discuss a setting in which coloring methods can outperform the algorithm from this paper.
@@ 1.3 Past work
To the best of our knowledge, Problems 1 and 2 have not been previously stated explicitly in the given generality. However, there is a range of past work that studies special cases of these problems. We categorize these works into the zero error case, where $\mathbf{A}=\mathbf{S}\circ\mathbf{A}$, and the nonzero error case, where $\mathbf{A}\neq\mathbf{S}\circ\mathbf{A}$. In addition, we discuss how Problems 1 and 2 differ from the standard sparse-recovery problem in compressed sensing.
@@ 1.3.1 Zero error
Some of the earliest work relating to Problems 1 and 2 seeks to recover Jacobian and Hessian matrices with a known sparsity pattern from matvecs [curtis_powell_reid_74, coleman_more_83, coleman_cai_86]. These methods make use of the fact that a graph coloring of a particular graph, induced by the sparsity pattern of $\mathbf{A}$, can be used to obtain a set of query vectors which are sufficient to exactly recover $\mathbf{A}$. In many (but not all) cases, exact recovery is possible with $s$ queries. Coloring methods have also influenced many algorithms for the nonzero error setting. We compare our results to such coloring-based methods in Section 4, arguing that Algorithm 1 always performs better in the zero-error setting, since it always requires just $s$ queries.
More generally, [halikias_townsend_23] studies the matvec query complexity of exact recovery of a wide range of linearly parameterized matrix families, proving matching upper and lower-bounds on the number of queries required. In particular, it is shown that recovering a diagonal matrix requires one query, recovering a block-diagonal matrix with $s\times s$ blocks requires $s$ queries, and recovering a tridiagonal matrix requires $3$ queries. Recovering a general $n\times d$ matrix is shown to require $d$ queries. With probability one, Algorithm 1 matches these lower bounds (see Proposition 1).
@@ 1.3.2 Nonzero error
$\mathbf{d}_{m}=\Big{[}\sum_{j=1}^{m}\mathbf{r}_{j}\circ(\mathbf{A}\mathbf{r}_{j})\Big{]}\ooalign{\hss$\circ$\hss\cr$\div$}\Big{[}\sum_{j=1}^{m}\mathbf{r}_{j}\circ\mathbf{r}_{j}\Big{]},$ (3)
@@
where “ $\circ$ $\div$ ” indicates entrywise division and the entries of the vectors $\mathbf{r}_{j}$ are all independent random variables with mean zero and variance one. A number of analyses of this estimator have been given for various distributions [bekas_kokiopoulou_saad_07, tang_saad_11, baston_nakatsukasa_22, hallman_ipsen_saibaba_23, dharangutte_musco_23]. In particular, [baston_nakatsukasa_22, dharangutte_musco_23] give error bounds for Problem 2, showing it suffices to set $m=O(1/\varepsilon)$, matching Theorem 1 up to constant factors. In fact, when $\mathbf{S}=\mathbf{I}$ our Algorithm 1 is equivalent to Equation 3 if the query vectors $\mathbf{r}_{i}$ are Gaussian. We detail this connection in Remark 3.
@@
Past work has also studied Problem 1 with the goal of approximating a potentially non-sparse matrix by a sparse matrix. For instance, there is a long line of work on approximating matrix functions $\mathbf{A}=f(\mathbf{H})$ from matvecs. If $\mathbf{H}$ is sparse, then the entries of $\mathbf{A}=f(\mathbf{H})$ decay exponentially away from the nonzero entries of $\mathbf{H}$ under mild assumptions on $f(x)$ [demko_moss_smith_84, benzi_razouk_08, benzi_boito_razouk_13]. As such, it is reasonable to approximate $\mathbf{A}$ with a sparse matrix of a sparsity similar to $\mathbf{H}$. This observation has been used in matrix approximation algorithms [tang_saad_11, stathopoulos_laeuchli_orginos_13, frommer_schimmel_schweitzer_21, park_nakatsukasa_23]. Broadly speaking, these algorithms aim to combine the coloring methods described above with the estimator $\mathbf{d}_{m}$ described in Equation 3. For banded matrices, one can use the existing analyses of $\mathbf{d}_{m}$ to analyze the performance of these methods, showing that they solve Problem 2 to accuracy ${\varepsilon}$ using $O(s/\varepsilon)$ matrix-vector queries. We include a note on this in Section 4.1, as we were unable to find such an analysis in the literature. Theorem 1 shows that Algorithm 1 matches this bound for arbitrary sparsity patterns.
@@
More recently, motivated by the field of partial differential equation (PDE) learning, there has been widespread interest in learning the solution operators of PDEs from input-output data of forcing terms and solutions, analogous to matrix-vector products [schafer_katzfuss_owhadi_21, schafer_owhadi_21, boulle_halikias_townsend_23, karniadakis_kevrekidis_lu_perdikaris_wang_yang_21]. The method in [schafer_owhadi_21] obtains a fixed-sparsity approximation to the sparse Cholesky factorization of the solution operator by coloring, which is provably accurate for certain problems. This makes use of the fact that in certain settings a fixed-sparsity Cholesky factorization is accurate and can be efficiently computed [schafer_katzfuss_owhadi_21]. This is broadly related to the (factorized) sparse approximate inverse problem for obtaining preconditioners [benzi_tuma_99]. The method in [boulle_halikias_townsend_23] also derives a continuous analogue of a generalized coloring algorithm for targeting low-rank subblocks of hierarchical matrices [levitt_martinsson_22a], and then recovering these subblocks using the randomized SVD [halko_martinsson_tropp_11]. The final step of this algorithm reduces to the recovery of a block diagonal matrix.
@@ 1.3.3 The sparse recovery problem in compressed-sensing
It is important to contrast the aims and methods of this paper with the rich literature on compressed sensing and sparse recovery [eldar_kutyniok_12, foucart_13, etc.].
@@
Critically, the support of $\widetilde{\mathbf{a}}$ is not known ahead of time. In fact, if the support were specified, this would be a trivial problem; simply take $\mathbf{M}$ to have $s$ rows, each a standard basis vector corresponding to an entry of the support.
A number of past works [waters_sankaranarayanan_baraniuk_11, wimalajeewa_eldar_varshney_13, dasarathy_shah_bhaskar_nowak_15, etc.] have also studied a matrix version of this problem in which one aims to obtain an $sd$-sparse matrix $\widetilde{\mathbf{A}}$ for which
@@
using only bi-linear measurements of $\mathbf{A}$: $(\mathbf{U},\mathbf{V})\mapsto\mathbf{U}\mathbf{A}\mathbf{V}^{\mathsf{T}}$. Note that the matrix problem is actually equivalent to a restricted version of the vector problem. Indeed, if $\mathbf{a}=\operatorname{vec}(\mathbf{A})$ and $\mathbf{M}=\mathbf{U}\otimes\mathbf{V}$, then $\mathbf{M}\mathbf{a}=\operatorname{vec}(\mathbf{U}\mathbf{A}\mathbf{V}^{\mathsf{T}})$. Here $\operatorname{vec}(\cdot)$ forms a vector by stacking the columns of a matrix on top of one another and $\otimes$ is the Kronecker product. In the zero error setting it has been shown that if $\mathbf{A}$ is an $n\times n$ matrix with sufficiently distributed sparsity, one can use a convex program to stably recover $\mathbf{A}$ using $O(\sqrt{\operatorname{nnz}(\mathbf{A})\cdot\log n})$ queries on each side [dasarathy_shah_bhaskar_nowak_15].
Even for the general vector recovery problem, such algorithms necessarily have worse dependencies on $\varepsilon$ and $d$ than the bounds we prove for the algorithm described in the next section. In particular, algorithms solving the $\ell_{2}/\ell_{2}$ sparse recovery problem necessarily require $\Omega(s/\varepsilon^{2}+s\log(ds)/\varepsilon)$ linear measurements [price_woodruff_11]. In fact, even if the problem is relaxed, so that the output vector $\widetilde{\mathbf{a}}$ is allowed to be non-sparse, $\Omega(s\log(d/s)/\varepsilon)$ queries are required [price_woodruff_11]. In contrast, our upper-bound Theorem 1/Corollary 1 have no dependence on the dimension $d$.